the price of a cellular telephone is based on peak and nonpeak service. kelsey used 45 peak and 50 nonpeak minutesand was charged $27.75. that same moth ,mitch used 70 peak minutes and 30 nonpeak minutes for a total charge of $36. what are the rates per minute for peak and nonpeak time?
i got 45p+50n=27.75
70p+30n=36.00
______________
were do i go from here

Answer by scott8148(6628)

multiplying 1st eqn by -3/5 ___ -27p - 30n = -16.65

adding 2nd eqn ___ 43p = 19.35

dividing by 43 ___ p = .45

substituting ___ 70(.45) + 30n = 36.00 ___ 31.50 + 30n = 36.00 ___ 30n = 4.50 ___ n = .15